# On long-duration storage, weather uncertainty and limited foresight

### v1

@@ 2.2 Foresight and stagewise independence
The multi-stage stochastic program formulation of a capacity expansion model presented in the last section did not specify the number of stages $T$, or the stage lengths $H(t)$. Any investigation of the effect of limited foresight on renewable energy systems naturally begs the question how limited the foresight actually is. The model implies that the system planner has perfect foresight for the duration of each stage.
Figure 2: Autocorrelation functions of weather variables at monthly stage length
@@
In the face of weather uncertainty this is directly related to the system planner’s ability to produce or procure reliable weather forecasts. Currently, mid-latitude weather forecasts are reliable up ten days out. While meteorologists see room for improvements, some have found an intrinsic, that is theoretical, limit to be 14 days in advance [bauer_quiet_2015, selz_transition_2022]. Broader developments could eventually be predicted a month ahead, as a joint research program by the University of Reading, the UK Met Office and the European Center for Medium-Range Weather Forecasts suggests [noauthor_how_2024].
@@
Brown et al. [brown_price_2025] opt for a 96 hour forecast horizon. The literature on hydroelectric energy systems typically assumes stage lengths from a week up to a month [hole_capacity_2025, shapiro_risk_2013-1], though the forecast of hydrological inflows may be different to that of solar irradiation and wind speeds.
@@
We choose a rather long stage length of one month for two reasons. For one, as we explain below, a monthly stage length is consistent with the stagewise independence assumption we made in the previous section. For two, assuming a longer foresight horizon that is likely to be conservative in that it underestimates the effect of weather uncertainty on capacity and dispatch decisions in a CEM. Shorter horizons would tend to amplify the effects observed in Section 4.
@@
In the previous section, we assumed stagewise independence. That is, the distribution of $\xi_{t+1}$ is independent of the distribution of $\xi_{t}$. This assumption is crucial for model tractability as it allows to approximate a single function of expected cost-to-go. If the expectation were conditional on the realization of $\xi_{t}$, e.g. a particularly windy stage $t$ would make it more likely to encounter a particularly windy state in $t+1$, the problem would grow substantially or require additional simplifications [lohndorf_modeling_2019].
@@
We devise a simple test for the stagewise independence assumption. As detailed in Section SI.10, we compute the autocorrelation function for the deviations from the stage mean of all the input parameters entering $\xi$. Figure 2 shows the autocorrelation for each input parameter for twelve monthly lags. Most importantly, solar PV, wind capacity factors and heat demand show no significant autocorrelation. Heat pump coefficients show a weakly significant first lag. By contrast, hydrological input factors for run-of-river, reservoirs and open pumped hydro storage inflows show significant autocorrelation. In the case study below, we exclude reservoirs and open pumped hydro to concentrate on hydrogen cavern LDES. While this simple test possibly neglects higher-order or cross-dependencies, we conclude that stagewise independence seems like a reasonable assumption for a monthly stage length. By contrast, we show in Section SI.10 that weekly stages exhibit some first-order autocorrelation even for solar PV and wind capacity factors.
@@ 2.3 Storage target level
The Limited Foresight model has twelve monthly subproblems, representing one operational year. CEMs are used to find cost-optimal long-run equilibria. Without any additional restriction, the system planner has no incentive to retain energy in LDES beyond the final stage of a problem formulation with finite stages and would empty the storage greedily, which would be at odds with the notion of a steady-state long-run equilibrium in which the operational year is a representation of many. In deterministic settings this is usually ensured by introducing a circularity constraint which connects the last time step to the first in the storage balance. In the stochastic setting considered here, this cannot be guaranteed.
Instead, we introduce an additional state variable for long-duration storage $s$, $e_{s}^{ini}$. The system planner chooses the initial storage level along with generation and storage capacities in stage 00 determining the starting endowment in stage $1$. In period $T$, any deviation of the ending storage level from $e_{s}^{ini}$ is penalized at a very high price (value of lost load). The system planner faces the trade-off of setting a high starting endowment at the risk of paying penalties if the storage target level is not met by the end of the model horizon.
Such storage targets are not far-fetched. Similar targets have been set for natural gas storage facilities by the European Commission for several years now [european_commission_commission_2024]. However, as we will see below, it represents coarse approximation of the marginal value of storage beyond the last stage, which may bias dispatch and capacity decisions in an end-of-horizon effect [shapiro_periodical_2020].
Hole et al. [hole_capacity_2025] recently used an discounted infinite horizon formulation as a remedy for a capacity expansion model. The length of forward and backward passes of the SDDP algorithm becomes random [dowson_policy_2020]. At reasonable discount rates, which determine the probability of different lengths, the number of subproblems to solve grows significantly. Computational tractability is maintained by a very reduced representation of dispatch operations.
@@ 2.4 Long-duration storage bidding behavior
We derive the LDES bidding behavior in the Limited foresight model building on Brown et al. [brown_price_2025]. Each dispatch stage is a linear program, for which the Karush Kuhn Tucker (KKT) conditions give necessary and sufficient optimality condition.
@@
where $\lambda_{t,h}$ is the electricity price in period $h$, $\lambda_{s,t,h}^{MSV}$ is the MSV/water value of LDES $s$, $\eta_{s}^{f}$ is the discharge efficiency and $\underline{\mu}_{s,t,h}^{f}$ and $\bar{\mu}_{s,t,h}^{f}$ are the dual variables to the H2 turbine (LDES dispatch) capacity bounds $[0,H_{s}]$. Likewise, the storage charging condition is given by,
@@
where $\underline{\mu}_{s,t,h}^{e}$ and $\bar{\mu}_{s,t,h}^{e}$ are the dual variables to the storage level bounds $[0,E_{s}]$ In the last period of each stage, the condition is given by,
@@
The partial (sub-)derivative of the expected cost-to-go with respect to the LDES ending level at the ending level is simply the expected MSV at the beginning of the next stage conditional on the LDES ending level. We also clarify that the expectation is taken with respect to next stage’s weather and can express the MSV as,
@@
This recursive structure is at the heart of LDES bidding. The MSV reflects the consequences of over all possible weather realizations from $t$ onward. If in a particular stage $t^{\prime}>t$ and under a particular weather trajectory the storage runs empty in period $h$, $\underline{\mu}_{s,t^{\prime},h}$ turns positive reflecting the outside option, load-shedding or imports. Such events in possible weather realizations from $t+1$ onward feed back to $\lambda_{s,t,h}^{MSV}$ through the tree of possible weather realization and enter the expectation. Thereby, the MSV represents the conditional probability of a bad system state times the marginal costs associated with such a state. The probability is conditional on the current storage level and on the remaining dispatch horizon. A higher storage level reduces the probability of bad states occurring in the future and the MSV goes down. Likewise, having a low storage level in November with imminent Dunkelflaute periods leads to higher MSVs than having the same storage level in March, when the worst of winter is over and the system operator can expect sufficient VRE supply in the coming months.
SDDP builds outer approximations of the piece-wise linear $\mathcal{Q}-$functions and we can use the trained model to back out the marginal storage values as a function of the current month (stage) and storage level. We do so by evaluating the partial (sub-)derivatives (slope) of $\{\hat{\mathcal{Q}}_{t}\}_{t=1}^{T}$ with respect to the (initial) storage level over the domain of the storage level $[0,E_{s}]$ in 10 GWh steps. We refer to these functions as LDES bidding curves.
For the last month $T$, the expected MSV is equal to the penalty for any ending level in $T-1$ for which one cannot meet the storage target level and equal to zero for any level that can.
@@ 2.5 Representation of risk preferences and outside options
In the context of interannual weather variability, the aim of CEMs is often to find resilient or robust system designs that avoid load shedding or the extensive use of an expensive backup technology. In a deterministic setting one can simply exclude the possibility of energy balance violations or attach a sufficiently high price for them not to occur. The stochastic setting considered here requires a more careful consideration of societal risk preferences.
@@
Coherent risk measures, like conditional value-at-risk, have a dual representation [shapiro_lectures_2009], which is equivalent to an optimization of an expectation over a changed probability distribution [liu_risk_2020] that puts more weight on adverse outcomes. A more risk-averse system operator would attach a higher probability to those scenarios in which the storage runs empty and the LDEs dispatch would be more conservative. Changing the costs of bad system states instead has a similar effect in that it increases the expected marginal value of storage. Without further formalization, we simply take away that assuming the availability of backup options at different costs may have a secondary interpretation of varying the degree of risk aversion the system operator exerts when operating the LDES. Along the same lines, previous research on risk-averse CEMs has shown that capacity decisions are driven by the costs of extreme scenarios, also in the risk-neutral case [munoz_does_2017].
In the base scenario of the case study below, we assume a very high VOLL to enable a direct comparison of the limited foresight model with a perfect foresight model that aims to be resilient to weather variability.
@@ 3 Case study
Figure 3: System representation and sampling lattice: Panel A shows is a stylised representation of the considered system. Electricity generation is confined to renewable sources. In case of Spain and UK, existing nuclear capacities are also considered. Electrified heat demand is served by a single type of heat pump.Hydrogen is produced via electrolysis or imported by ship, either under long-term contracts or, depending on the scenario, via spot shipments. Hydrogen can be stored in pressurised steel tanks or caverns and combusted in turbines to generate electricity. Panel B is a partial representation of the scenario lattice representing sample space $\Xi$. Each node represents a historical weather month. A sample path $\xi_{i}$ can be any combination of monthly nodes, like the trajectory of darkblue nodes.
@@ 3.1 Model description
For both models, we use an adapted version of the open-source dispatch and capacity expansion model Dispatch and Investment Evaluation Tool with Endogenous Renewables (DIETER) [zerrahn_long-run_2017, schill_long-run_2018]. The model has been used in various studies to investigate flexibility in sector-coupled VRE-based energy systems, e.g. [zerrahn_economics_2018, schill_electricity_2020, roth_geographical_2023, roth_power_2024, gueret_impacts_2024]. DIETER is a linear cost-minimizing model that abstracts from discrete capacity decisions and economies of scale as well as non-convex operational constraints. The model is centered around the power sector with coupled hydrogen and heating sectors, see Figure 3, upper panel. We describe a simplified formal model in Section SI.5.
The model version used in this case study strikes a balance between capturing most of the relevant dynamics in a sector-coupled decarbonized energy system and maintaining computational tractability. As Section SI.7 shows, SDDP solves millions of monthly subproblems in order to identify an approximately optimal policy. We reduce the computational burden by averaging all time series to four-hourly time steps, which tends to have negligible effects on optimal capacities [schlachtberger_cost_2018]. We also solve separate problems for three single-node systems, Germany, Spain and the United Kingdom (UK), representing a balanced, a solar-dominated and a wind-dominated system respectively, as done in other studies concentrating on LDES and weather variability [brown_ultra-long-duration_2023] . We abstract from any electricity or hydrogen exchange with neighboring countries save for hydrogen imports at exogenous prices. The modeled target year is 2050 and only fossil-free generation technologies are admissible candidates in a static, one-shot capacity investment stage.
@@ 3.2 Input data
The candidate generation portfolio is comprised of solar PV, on- and offshore wind, bioenergy, run-of-river hydroelectric plants and nuclear capacity for the case of Spain and UK. Storage technologies in the power sector include lithium-ion batteries and closed pumped-hydro storage. We assume that these short-term storages are subject to a circularity constraint within each month. We abstract from other hydro-storage technologies with inflow. [^2] Reservoirs are another type of LDES. Their inclusion would consequently increase the state space of the SDDP model, as their storage level would also need to be tracked over the months. The stagewise independence test in Section 2.2 additionally suggests that there is some autoregressive structure in hydro inflows which would futher complicate adequate representation.
For the hydrogen (H2) sector, the model can invest in polymer electrolyte membrane (PEM) electrolysis, underground H2 storage in salt caverns and pressurized tank storage above ground as well as hydrogen gas turbines for reconversion to electricity. Additionally, the system operator can enter into long-term contracts (LTC) for H2 that deliver a constant amount in every time step subject to limited offtake flexibility of +/-10%.
@@
Almost all cost and technical parameters are taken from the Danish Energy Agency projections for 2050 [danish_energy_agency_technology_2024]. We provide more details in Table SI.3. Technical capacity bounds for solar PV and wind onshore are taken from Tröndle et al. [trondle_home-made_2019]. Wind offshore is constrained by the upper limit in the Ten Year Network Development Plan (TYNDP) 2024 of the European Network of Transmission System Operators for electricity and gas (ENTSO-E and ENTSO-G) [entso-e_and_entso-g_tyndp_2024] to account for other build-out constraints than just technical availability. As a lower bound for solar PV and wind we take currently installed capacities. Biomass, run-of-river and nuclear generation capacities is constrained to National Estimates for 2030 in the European Resource Adequacy Assessment (ERAA) 2021 [entso-e_european_2021]. We provide details in Table SI.4 and SI.5. Cavern storage potentials are taken from Caglayan et al. [caglayan_technical_2020].
For weather data, we use country-level data for solar PV, wind and run-of-river capacity factors from the Pan-European Climate Database [de_felice_entso-e_2022] and heat demand and heat pump coefficients of performance (COP) from Göke [goke_pecd_2024]. We obtain hourly time series for 1982-2017. We average the data to four-hourly periods and define summer-to-summer weather years from July to June. For each calendar month $t$, we compile a sample space $\Xi_{t}$ that contains 35 historical samples from that month, i.e. $\Xi_{1}$ contains data for 35 Julys from 1982/83 to 2016/17, $\Xi_{2}$ contains 35 Augusts and so on. Each sample comprises 168 to 186 time steps. This way, the sampling process preserves the inherent seasonality of the weather data. Figure 3, lower panel, illustrates the stagewise independent sampling process. Any combination of stratified weather months is admissible resulting in a total sampling space $\Xi=\Xi_{1}\times\cdots\times\Xi_{12}$ of $|\Xi|=35^{12}\approx 3.38\times 10^{18}$ possible synthetic weather years.

### v2

@@ 2.2 Foresight and stagewise independence
The multi-stage stochastic program formulation of a capacity expansion model presented in the last section did not specify the number of stages $T$, or the stage lengths $H(t)$. Any investigation of the effect of limited foresight on renewable energy systems naturally begs the question of how limited foresight actually is. The model implies that the system planner has perfect foresight for the duration of each stage.
Figure 2: Autocorrelation functions of weather variables at monthly stage length. Monthly autocorrelation for solar photovoltaic, wind onshore, wind offshore, run of river, hydro reservoirs, pumped hydro storage (PHS) with inflows, residential and commercial aggregate heat demand and heat pump efficiencies.
@@
In the face of weather uncertainty, this is directly related to the system planner’s ability to produce or procure reliable weather forecasts. Currently, mid-latitude weather forecasts are reliable up to ten days out. While meteorologists see room for improvements, some have found an intrinsic, that is, theoretical limit to be 14 days in advance [bauer_quiet_2015, selz_transition_2022]. Broader developments could eventually be predicted a month ahead, as a joint research program by the University of Reading, the UK Met Office and the European Center for Medium-Range Weather Forecasts suggests [national_centre_for_atmospheric_science_how_2024].
@@
Brown et al. [brown_price_2025] opt for a 96-hour forecast horizon. The literature on hydroelectric energy systems typically assumes stage lengths from a week up to a month [hole_capacity_2025, shapiro_risk_2013]. Yet, the forecast for hydrological inflows may be different from that of solar irradiation and wind speeds.
@@
We choose a rather long stage length of one month for two reasons. For one, as we explain below, a monthly stage length is consistent with the stagewise independence assumption we made in the previous section. For another, assuming a longer foresight horizon is likely to be conservative in that it underestimates the effect of weather uncertainty on capacity and dispatch decisions in a CEM. Shorter horizons would tend to amplify the effects observed in Section 4.
@@
In the previous section, we assumed stagewise independence. That is, the distribution of $\xi_{t+1}$ is independent of the distribution of $\xi_{t}$. This assumption is crucial for model tractability as it allows to approximate a single function of expected cost-to-go. If, for example, a particularly windy stage $t$ were to make it more likely to encounter a particularly windy state in $t+1$, the problem would grow substantially in size or require additional simplifications [lohndorf_modeling_2019].
@@
We devise a simple test for the stagewise independence assumption. As detailed in Section .10, we compute the autocorrelation function for the deviations from the stage mean of all the input parameters entering $\xi$. Figure 2 shows the autocorrelation for each input parameter for twelve monthly lags. Most importantly, solar PV, wind capacity factors and heat demand show no significant autocorrelation. Heat pump coefficients show a weakly significant first lag. By contrast, hydrological input factors for run-of-river, reservoirs and open pumped hydro storage inflows show significant autocorrelation. In the case study below, we exclude reservoirs and open pumped hydro to concentrate on hydrogen cavern LDES. While this simple test possibly neglects higher-order or cross-dependencies, we conclude that stagewise independence seems like a reasonable assumption for a monthly stage length. By contrast, we show in Section .10 that weekly stages exhibit some first-order autocorrelation even for solar PV and wind capacity factors.
@@ 2.3 Storage target level
The Limited Foresight model has twelve monthly subproblems, representing one operational year. CEMs are commonly used to find cost-optimal long-run equilibria. Without any additional restriction, the system planner has no incentive to retain energy in LDES beyond the final stage of a problem formulation with finite stages and would empty the storage greedily. This would be at odds with the notion of a steady-state long-run equilibrium, in which the operational year is a representation of many. In deterministic settings, this is usually ensured by introducing a circularity constraint which connects the last time step to the first in the storage balance. In the stochastic setting considered here, this cannot be guaranteed.
Instead, we introduce an additional state variable for long-duration storage $s$, $e_{s}^{ini}$. The system planner chooses the initial storage level along with generation and storage capacities in stage 00, which determines the starting endowment at the start of stage $1$. In period $T$, any deviation of the ending storage level from $e_{s}^{ini}$ is penalized at a very high price (value of lost load). The system planner faces the trade-off of setting a high starting endowment at the risk of paying penalties if the storage target level is not met by the end of the model horizon.
Such storage targets are not far-fetched. Similar targets have been set for natural gas storage facilities by the European Commission for several years now [european_commission_commission_2024]. However, as we will see below, it represents a coarse approximation of the marginal value of storage beyond the last stage, which may bias dispatch and capacity decisions in an end-of-horizon effect [shapiro_periodical_2020].
Hole et al. [hole_capacity_2025] recently used a discounted infinite horizon to eliminate such end-of-horizon effects in a capacity expansion model. The length of forward and backward passes of the SDDP algorithm becomes random [dowson_policy_2020]. At reasonable discount rates, which determine the expected pass lengths, the number of subproblems to solve grows significantly. Computational tractability is maintained by a very reduced representation of dispatch operations.
@@ 2.4 Long-duration storage bidding behavior
We derive the LDES bidding behavior in the Limited foresight model, building on Brown et al. [brown_price_2025]. Each dispatch stage is a linear program, for which the Karush-Kuhn-Tucker (KKT) conditions give necessary and sufficient optimality conditions.
@@
where $\lambda_{t,h}$ is the electricity price in period $h$, $\lambda_{s,t,h}^{MSV}$ is the MSV of LDES $s$, $\eta_{s}^{f}$ is the discharge efficiency and $\underline{\mu}_{s,t,h}^{f}$, and $\bar{\mu}_{s,t,h}^{f}$ are the dual variables to the hydrogen turbine (LDES dispatch) capacity bounds $[0,H_{s}]$. Likewise, the storage charging condition is given by,
@@
where $\underline{\mu}_{s,t,h}^{e}$ and $\bar{\mu}_{s,t,h}^{e}$ are the dual variables to the storage level bounds $[0,E_{s}]$ In the last period $H(t)$ of each stage, the condition is given by,
@@
The partial (sub-)derivative of the expected cost-to-go with respect to the LDES ending level at the current ending level is simply the expected MSV at the beginning of the next stage, conditional on the LDES ending level $e_{s,t,H}$. We also clarify that the expectation is taken with respect to the next stage’s weather vector $\xi_{t+1}$ and can express the MSV in stage $t$ and period $h$ as,
@@
This recursive structure is at the heart of LDES bidding under uncertainty. The MSV reflects the consequences of all possible weather realizations from $t$ onward. If in a stage $t^{\prime}>t$ under a particular weather trajectory, the storage runs empty in period $h$, $\underline{\mu}_{s,t^{\prime},h}$ turns positive, reflecting the cost of an outside option, e.g. load-shedding or energy imports. Such events in specific weather realizations from $t+1$ onward feed back to $t$ through the nested tree of possible weather realizations and enter the expectation in the expression for $\lambda_{s,t,h}^{MSV}$. Thereby, the MSV represents the conditional probability of a bad system state times the marginal costs associated with such a state. The probability is conditional on the current storage level and on the remaining dispatch horizon. A higher storage level reduces the probability of bad states occurring in the future and the MSV goes down. Likewise, having a low storage level in November with imminent Dunkelflaute periods leads to higher MSVs than having the same storage level in March, when the worst of winter is over and the system operator can expect sufficient VRE supply in the coming months.
SDDP builds outer approximations of the piece-wise linear $\mathcal{Q}-$functions, and the trained model can be used to back out the marginal storage values as functions of the current month (stage) and storage level. We do so by evaluating the partial (sub-)derivatives (slope) of $\{\hat{\mathcal{Q}}_{t}\}_{t=1}^{T}$ with respect to the stage’s LDES ending level over the domain of the storage level $[0,E_{s}]$ in 10 GWh steps. We refer to these functions as LDES bidding curves.
For the last month $T$, the expected MSV is equal to the penalty for any ending level below the initial storage level $e^{ini}_{s}$ and zero above.
@@ 2.5 Representation of risk preferences and outside options
In the context of interannual weather variability, the aim of CEMs is often to find resilient or robust system designs that avoid load shedding or the extensive use of an expensive backup technology. In a deterministic setting, one can simply exclude the possibility of energy balance violations or attach a sufficiently high price for them not to occur. The stochastic setting considered here requires a more careful consideration of societal risk preferences.
@@
Coherent risk measures, like conditional value-at-risk, have a dual representation [shapiro_lectures_2009], which is equivalent to an optimization of an expectation over a changed probability distribution [liu_risk_2020] that puts more weight on adverse outcomes. A more risk-averse system operator would attach a higher probability to those scenarios in which the storage runs empty and the LDES dispatch would be more conservative. It should be noted that increasing the costs of bad system states instead of attaching a higher probability has a similar effect in that it increases the expected marginal value of storage. Without further formalization, we simply take away that varying the availability and costs of backup options may have a secondary interpretation of varying the degree of risk aversion the system operator exerts when operating the LDES.
In the case study below, we vary the availability of outside options. The base scenario assumes a very high value of lost load and no alternative backup to enable a direct comparison of the limited foresight model with a perfect foresight model that aims to be resilient to weather variability.
@@ 3 Case study
Figure 3: System representation and sampling lattice: Panel A shows a stylized representation of the considered system. Electricity generation is confined to renewable sources. In the case of Spain and the UK, existing nuclear capacities are also considered. Electrified heat demand is served by a single type of heat pump. Hydrogen is produced via electrolysis or imported by ship, either under long-term contracts or, depending on the scenario, via spot shipments. Hydrogen can be stored in pressurized steel tanks or caverns and combusted in turbines to generate electricity. Panel B is a partial representation of the scenario lattice representing sample space $\Xi$. Each node represents a historical weather month. A sample path $\xi_{i}$ can be any combination of monthly nodes, like the example trajectory of darkblue nodes.
@@ 3.1 Model description
For both models, we use an adapted version of the open-source dispatch and capacity expansion model Dispatch and Investment Evaluation Tool with Endogenous Renewables (DIETER) [zerrahn_long-run_2017, schill_long-run_2018]. The model has been used in various studies to investigate flexibility in sector-coupled VRE-based energy systems [e.g. zerrahn_economics_2018, schill_electricity_2020, roth_geographical_2023, roth_power_2024, gueret_impacts_2024]. DIETER is a linear cost-minimizing model that abstracts from discrete capacity decisions and economies of scale as well as non-convex operational constraints. The model is centered around the power sector with coupled hydrogen and heating sectors, see Figure 3, panel A. We describe a simplified formal model in Section SI.5.
The model version used in this case study strikes a balance between capturing most of the relevant dynamics in a sector-coupled decarbonized energy system and maintaining computational tractability. As Section SI.7 shows, the SDDP algorithm solves millions of monthly subproblems in order to identify an approximately optimal policy. We reduce the computational burden by averaging all time series to four-hourly time steps, which tends to have negligible effects on optimal capacities [schlachtberger_cost_2018]. We solve separate problems for three single-node systems. Germany, Spain and the United Kingdom (UK) represent a balanced, a solar-dominated and a wind-dominated system respectively. These countries have also been used in other studies concentrating on LDES and weather variability [brown_ultra-long-duration_2023]. We abstract from any electricity or hydrogen exchange with neighboring countries, save for hydrogen imports at exogenous prices. The modeled target year is 2050, and only zero-emission generation technologies are admissible candidates in a static, one-shot capacity investment stage.
@@ 3.2 Input data
The candidate generation portfolio is comprised of solar PV, on- and offshore wind, bioenergy, run-of-river hydroelectric plants and nuclear capacity for the case of Spain and UK. Storage technologies in the power sector include lithium-ion batteries and closed pumped-hydro storage. We assume that these short-term storages are subject to a circularity constraint within each month. We abstract from other hydro-storage technologies with inflow. [^2] Reservoirs would require a similar treatment as LDES. Their inclusion would consequently increase the state space of the SDDP model, as their storage level would also need to be tracked over the months. The stagewise independence test in Section 2.2 additionally suggests that there is some autoregressive structure in hydro inflows, which would further complicate an adequate representation.
For the hydrogen (H2) sector, the model can invest in polymer electrolyte membrane (PEM) electrolysis, underground H2 storage in salt caverns and pressurized tank storage above ground as well as hydrogen gas turbines for reconversion to electricity. Additionally, the system operator can enter into long-term contracts (LTC) for H2 that deliver a constant amount in every time step, subject to limited off-take flexibility of +/-10%.
@@
Almost all cost and technical parameters are taken from the Danish Energy Agency projections for 2050 [danish_energy_agency_technology_2024]. We provide more details in Table SI.3. Technical capacity bounds for solar PV and wind onshore are taken from Tröndle et al. [trondle_home-made_2019]. Wind offshore is constrained by the upper limit in the Ten Year Network Development Plan (TYNDP) 2024 of the European Network of Transmission System Operators for electricity and gas (ENTSO-E and ENTSO-G) [entso-e_and_entso-g_tyndp_2024] to account for other build-out constraints than just technical availability. As a lower bound for solar PV and wind we take the currently installed capacities. Biomass, run-of-river and nuclear generation capacities are constrained to National Estimates for 2030 in the European Resource Adequacy Assessment (ERAA) 2021 [entso-e_european_2021]. We provide details in Tables SI.4 and SI.5. Cavern storage potentials are taken from Caglayan et al. [caglayan_technical_2020].
For weather data, we use country-level time series for solar PV, wind and run-of-river capacity factors from the Pan-European Climate Database [de_felice_entso-e_2022] and heat demand and heat pump coefficients of performance (COP) data from Göke [goke_pecd_2024]. We obtain hourly time series for 1982-2017. We average the data to four-hourly periods and define summer-to-summer weather years from July to June. For each calendar month $t$, we compile a sample space $\Xi_{t}$ that contains 35 historical samples from that month, i.e. $\Xi_{1}$ contains data for 35 Julys from 1982/83 to 2016/17, $\Xi_{2}$ contains 35 Augusts and so on. Each sample comprises 168 to 186 time steps. This way, the sampling process preserves the inherent seasonality of the weather data. Figure 3, panel B, illustrates the stagewise independent sampling process. Any combination of stratified weather months is admissible, resulting in a total sampling space $\Xi=\Xi_{1}\times\cdots\times\Xi_{12}$ of $|\Xi|=35^{12}\approx 3.38\times 10^{18}$ possible synthetic weather years.